The attendence at home basketball games has ranged between 820 and 1540 spectators. Write an absolute value inequality descibing the possible attendance at home basketball games.
Answer by stanbon(75887)

The attendence at home basketball games has ranged between 820 and 1540 spectators. Write an absolute value inequality descibing the possible attendance at home basketball games.
------------
Find the average attendance figure: (820+1540)/2 = 1180
------------
Find the difference between average and max (or min). 1540-1180 = 360
----------------------------
Equation:
|x-1180| <= 360
x is the possible attendance at home basketball games.
======================
Cheers,
Stan H.